OpenStudy (anonymous):

A member of a car-sharing program can use a car for 6$ per hour and $.50 per mile. The member uses the car for one day and is charged $44. This situation is given by. 6t+0.5d=44. Where t is the time in hours the car is used and d is the distance in miles the car is driven. Give three examples of the number of hours the member could have used the car and the number of miles the member could have driven the car.

OpenStudy (anonymous):

6t+0.5d=44 Made a boo boo last post but corrected it :-D This is in standard form and all you have to do is evaluate. So plug some number into t and find d. 6(1)+0.5d=44 6 + 0.5d = 44 0.5d = 44-6 0.5d/.5 = 38/.5 d= 76 miles So for one hour he could of went 76 miles 6(2)+0.5d=44 12 + 0.5d = 44 0.5d = 44 - 12 0.5d = 32 0.5d/.5 = 32/.5 d = 64 He could of went 64 miles in 2 hours